If you could help me with this problem it would be greatly appricated! Thank you. Problem below:
Solve -6m-6+8m=-5+2m-1. Tell whether the equation has infinitely many solutions or no solutions.

Answer by dolly(163)


-6m-6+8m=-5+2m-1
On simplifying both the sides we get,
2m - 6 = 2m - 6
As both the sides are equal, the given equation has no solution at all
Good Luck!!!